Fermi National Laboratory
 

Operation Summary for October 23, 2009 - October 26, 2009

 

The Integrated Luminosity for the period from 10/19/09 to 10/26/09 was 56.26 pb-1.

Friday Oct. 23

The day shift began with Operations monitoring the Tevatron (TeV) store 7286 (68.81E30) and the Recycler stash (363.3E10), with the Antiproton Source stacking (10.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7286 at 9:07 AM. Operations prepared CDF and D0 for an access that lasted until 1 PM. During this access FESS personnel cleaned the TeV sector F0 pond water strainer. Operators checked the orbits and then put the TeV into HEP shot setup. Linac personnel worked on a RF station (LRF1) at the end of the shift.

During the evening shift, Operations established HEP store 7289 at 4:32 PM with an average initial luminosity (AIL) of 275.67E30. Operations tuned the MI beam to correct beam positions.

Saturday Oct. 24

The midnight shift began with Operations monitoring store 7289 (113.32E30) and the stash (303.4E10), with Pbar stacking (28.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported short downtimes due to the following trips: MI RF South Anode, MIRF4, a NuMI power supply (E:V118), and a Linac Klystron RF station (KRF4).

During the day shift, Operations terminated store 7289 and put the TeV into shot setup at 8:31 AM. Operations established store 7290 at 10:32 AM with an AIL of 274.71E30.

During the evening shift, LRF2 tripped off; it reset. Operations had to reboot the MI vacuum system front end.

Sunday Oct. 25

The midnight shift began with Operations monitoring store 7290 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7290 and put the TeV into shot setup at 1:41 AM. Operations reported that shot setup was extended due to problems with the H- Source vacuum and MIRF4. Operations established store 7291 at 4:16 AM with an AIL of 251.14E30. Operations reported short downtimes due to the following trips: Booster vacuum pump (BIPRF17), BRF18, and LRF2.

During the day shift, Operations reported beam permit trips for the P1 transfer line due to high current on a power supply (V701).

During the evening shift, LRF3 tripped off; it reset. P1 transfer line trips continued due to V701 high current and also a quadrupole power supply (I:Q703). The NuMI beam permit tripped off due to low current on a power supply (V118). Operations terminated store 7291, put the TeV into shot setup, and established store 7292 at 10:53 PM with an AIL of 278.97E30.

Monday Oct. 26

The midnight shift began with Operations monitoring store 7292 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported another alarm for the TeV sector F0 pond water pressure; this will be checked during the day shift. Q703 tripped off the P1 line beam permit; it reset. Operations reported problems with the Pbar autotune program; a reboot fixed the problem. At 5:34 AM, BRF18 began tripping off; it reset.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster and Pbar studies. She will also allow TeV personnel to change the D0 collision point.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/26/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory